### v1

# Shell quenching in nuclear charge radii based on Monte Carlo dropout Bayesian neural network

Abstract
Charge radii can be generally used to encode the information about various fine structures of finite nuclei. In this work, a constructed Bayesian neural network based on Monte Carlo dropout approach is proposed to accurately describe the charge radii of nuclei with proton number $Z\geq 20$ and mass number $A\geq 40$. More motivated underlying mechanisms are incorporated into this combined model except for the basic building blocks with the specific proton and neutron numbers, which naturally contains the pairing effect, isospin effect, shell closure effect associated to the Casten factor $P$, valence neutrons, valence protons, quadrupole deformation $\beta_{20}$, high order hexadecapole deformation $\beta_{40}$, and the local “abnormal” shape staggering effect of 181,183,185Hg. To avoid the over-fitting puzzle along $Z=28$, $50$, and $82$ isotopic chains, the modified Casten factor $P^{*}$ is tentatively introduced into the input structure parameter sets. We have successfully demonstrated the ability of the Monte Carlo dropout Bayesian neural network (MC-dropout BNN) model to significantly increase the accuracy in predicting the nuclear charge radii. The standard root-mean-square deviation falls into $0.0084$ fm for the training data and $0.0124$ fm for the validation data with the modified Casten factor $P^{*}$ input. Meanwhile, the shell closure effect of nuclear charge radii can be reproduced remarkably well.
I Introduction
Finite nuclei, which are characterized as the strongly correlated many-body systems, provide a natural probe to recognize the fundamental interactions and forces. As directly measured quantities in atomic nuclei, charge radii can be generally used to reflect the proton density distributions and encode the information about various fine structure phenomena, such as the emergence of shell closure effects Angeli_2015 ; PhysRevC.100.044310 ; GarciaRuiz:2019cog ; PhysRevC.104.064313 ; PhysRevLett.129.142502 and the odd-even staggering (OES) behavior GarciaRuiz:2016ohj ; Marsh:2018wxs ; Miller2019 ; PhysRevLett.127.192501 . Furthermore, available description of proton density distributions can also be used to pin down the components of isospin interactions in the equation of state (EoS) of asymmetric nuclear matter PhysRevC.88.011301 ; PhysRevLett.119.122502 ; PhysRevResearch.2.022035 ; PhysRevLett.127.182503 ; nuclscitech34.119 ; PhysRevLett.132.162502 . This means that more data about nuclear charge radii should be urgently required with the highly accurate precision. Remarkable achievements have been obtained in detecting charge radii over the past decades. Especially for nuclei with extreme neutron-to-proton ratios, laser spectroscopy method can be considered as an available access to measure the nuclear size CAMPBELL2016127 ; YANG2023104005 .
The undertaken efforts have been devoted to featuring the normal and abnormal variations in nuclear charge radii along a long isotopic family, which includes the empirical formulas Zhang2002EPJA285 ; PhysRevC.87.024310 ; Piekarewicz2010 ; PhysRevC.94.064315 , relativistic mean field theory Geng:2003pk ; PhysRevC.102.024307 ; PhysRevC.109.064302 , non-relativistic Skyrme Hartree-Fock-Bogoliubov (HFB) approach PhysRevLett.102.242501 ; PhysRevC.82.035804 , and the Fayans energy density functional PhysRevC.95.064328 ; PhysRevLett.122.192502 , etc. Recently, machine learning methods have been made greatly applications in the course of nuclear physics  PhysRevC.98.034318 ; Bedaque2021 ; Paquet_2024 . Particularly, a well-known approach based on Bayesian neural network (BNN) has been widely and successfully employed to describe various nuclear quantities from finite nuclei to infinite nuclear matter, which includes the nuclear mass PhysRevC.93.014311 ; PhysRevC.96.044308 ; NIU201848 ; PhysRevC.100.054311 ; Kejzlar_2020 ; PhysRevC.101.051301 ; PhysRevC.104.014315 ; PhysRevC.106.014305 ; PhysRevC.106.L021303 ; ZHANG2024122820 , $\alpha$-decay properties Rodríguez_2019 ; PhysRevC.107.014310 ; PhysRevC.108.014326 ; PhysRevC.110.024319 and $\alpha$-clustering structure PhysRevC.104.044902 , $\beta$-decay properties PhysRevC.99.064307 ; PhysRevC.106.024306 ; PhysRevC.107.034316 or double-$\beta$-decay models PhysRevD.96.053001 ; PhysRevD.104.055040 ; PhysRevC.94.024603 , $\gamma$-ray emission PhysRevC.105.045801 , nuclear fission Schunck_2020 , low-lying excited phenomena PhysRevC.104.034317 ; WANG2022137154 , landscape of nuclide chart PhysRevC.101.044307 ; PhysRevLett.122.062502 , neutron skin thickness PhysRevC.107.045802 ; PhysRevC.108.024317 , the yields of production cross sections PhysRevC.107.064909 ; PhysRevC.108.044606 ; YingHuaDang , and nuclear charge radii Utama_2016 ; PhysRevC.101.014304 ; PhysRevC.102.054323 ; PhysRevC.105.014308 ; DONG2023137726 , etc. Besides, the Bayesian inference has a guidance to gain deeper insights into the understanding of fundamental interactions PhysRevC.105.044305 ; Wesolowski_2016 ; PhysRevC.97.025805 ; YANG2020135540 ; PhysRevC.103.025807 ; XU2020135820 ; PhysRevC.103.065804 ; PhysRevC.105.014004 , significantly the bulk properties of nuclear matter PhysRevC.96.065805 ; PhysRevLett.125.202702 ; PhysRevC.103.064323 ; PhysRevD.103.063036 ; Xie_2021 ; PhysRevC.105.015806 ; PhysRevC.106.044305 ; PhysRevD.106.043024 ; panglonggang ; PhysRevC.107.045803 ; PhysRevC.108.034908 ; QIU2024138435 ; RADUTA2024138696 ; PhysRevC.107.055803 ; PhysRevC.110.035805 ; PhysRevD.109.123038 . It should be mentioned that Bayesian neural networks can cover the quantitative uncertainties associated with these various physical systems Higdon2015 ; PhysRevC.96.024003 ; PhysRevC.98.024305 ; PhysRevLett.119.252501 ; HU2019134982 ; PhysRevC.101.055803 ; PhysRevC.102.024616 ; PhysRevC.102.054315 ; Lovell_2020 ; PhysRevC.104.054324 ; PhysRevC.106.024607 ; PhysRevC.106.054001 ; PhysRevC.108.024321 ; PhysRevC.109.054301 ; XinyuWang2024 . This encourages us to perform the proceeding researches in exploring the systematical evolution of nuclear charge radii through Bayesian neural network based approach.
Accurate description of nuclear charge radii can be applied to refine the local variations in nuclear size. In Ref. Utama_2016 , it demonstrates that the BNN can increase the prediction ability of nuclear charge radii by training the residuals between the theoretical predictions and experimental data. This leads to the root-mean-square (rms) deviation with being around $0.02$ fm. The improved BNN which considers the pairing and shell closure effects gives the rms deviation $0.015$ fm in calibrating the charge radii data of nuclei with mass number $A\geq 40$ and proton number $Z\geq 20$ PhysRevC.105.014308 . As is well known, the larger odd-even staggering in charge radii can be profoundly observed in the neutron-deficient isotopes around the proton number $Z=82$ isotopic chains Marsh:2018wxs ; PhysRevLett.127.192501 ; PhysRevC.95.044324 ; ANSELMENT1986471 ; PhysRevC.104.024328 ; PhysRevLett.126.032502 . This is attributed to the underlying mechanism of the abrupt shape-phase transition. By considering the abnormal shape staggering effect, the standard rms deviation falls into $0.014$ fm for both the trained and validated charge radii data through the refined Bayesian neural network DONG2023137726 . In addition, different kinds of machine learning methods, such as the naive Bayesian probability classifier PhysRevC.101.014304 and artificial neural networks PhysRevC.102.054323 ; Akkoyun_2013 ; PhysRevC.108.034315 , can be used to explore the systematical evolution of nuclear charge radii as well.
As mentioned above, the Bayesian neural networks can capture the considered error quantification through comparing the experimental data and theoretical simulations. Moreover, this probabilistic model is more computational saving with respect to the generally used energy density functionals (EDFs). As mentioned in Ref. Zhang10662945 , the aleatoric and epistemic uncertainties should be explicitly considered in a given Bayesian neural network. This means that more complex input structures should be taken into account in the simulating process. In other words, more underlying physical mechanisms should be captured as much as possible in the calibrated protocol. To further reduce the quantified uncertainty in validating the nuclear charge radii, both of the quadrupole deformation $\beta_{20}$ and higher moment deformation $\beta_{40}$ are incorporated into the training process in this work. The Casten factor $P$, which is characterized as the correlation between the valence neutrons $\nu_{n}$ and the valence protons $\nu_{p}$, is associated to the shell closure effect PhysRevLett.58.658 ; Angeli_1991 ; Casten_1996 . In our calculations, the Casten factor $P$, valence neutrons $\nu_{n}$, and valence protons $\nu_{p}$ are simultaneously considered as the input parameter sets.
Actually, Casten factor $P$ plays an indispensable role in the calibrated protocol PhysRevC.105.014308 ; DONG2023137726 ; Sheng:2015poa . For the calculation of the Casten factor $P$, the reference neutron and proton magic numbers are taken as $Z=2$, $6$, $14$, $28$, $50$, $82$, $(114)$ and $N=2$, $8$, $14$, $28$, $50$, $82$, $126$, $(184)$ PhysRevC.88.011301 ; Angeli_1991 ; Sheng:2015poa . However, the traditional Casten factor, which is defined as $P=\nu_{n}\nu_{p}/(\nu_{n}+\nu_{p})$ is invalid along $Z=28$, $50$, $82$ isotopic chains due to the fully filled proton magic shells. In order to avoid the aforementioned problem, the shell closure effect has been taken into account through the modified Casten factor $P^{*}$. In this work, the Monte Carlo dropout Bayesian neural network (MC-dropout BNN) approach is built appropriately to describe the charge radii of nuclei with proton number $Z\geq 20$ and mass number $A\geq 40$. The input structures include the proton number ($Z$), mass number ($A$), the pairing effect ($\delta$), isospin effect ($I^{2}$), shell closure effect associated to the Casten factor $P^{*}$, valence neutrons $\nu_{n}$, valence protons $\nu_{p}$, quadrupole deformation $\beta_{20}$, high order hexadecapole deformation $\beta_{40}$, and the local “abnormal” shape staggering effect of 181,183,185Hg ($LI$). The results obtained by incorporating the traditional Casten factor $P$ into the input structures are also shown for comparison.
The structure of the paper is the following. In Sec. II, the theoretical framework about the MC-dropout BNN approach is briefly presented, which includes the engineered features and the training and validation data sets in the nuclear charge radii. In Sec. III, the numerical results and discussion are provided. Finally, a summary is given in Sec. IV.
II Theoretical formalism
The Bayesian neural network (BNN), which is characterized as the combination of an artificial neural network (ANN) and the Bayesian statistical theory, provides a potential approach to comprehend the physical quantities with the associated uncertainty. However, as demonstrated in Ref. JMLRv15srivastava14a , the over-fitting puzzle can be encountered in the traditional neural network. Then the proposed dropout approach can be used to avoid over-fitting problem of the neural network parameters through the regularization technique. Meanwhile, the constructed Monte Carlo dropout (MC-dropout) method is being equivalent to variational inference Bayesian learning method in evaluating a neural network Gal2016 . Therefore the Monte Carlo dropout Bayesian neural network (MC-dropout BNN) model has been built based on these motivations. This approach has been successfully used to analyze various physical quantities, such as the spectral function PhysRevResearch.4.043082 , mass distributions of the induced fission Huo2023 , and electron-carbon scattering data PhysRevC.110.025501 .
The dropout architecture with a single hidden layer can be recalled as follows,
$\hat{y}=\sigma(xz_{1}W_{1}+b)z_{2}W_{2},$ (1)
where $\sigma$ depicts the nonlinear activation function, $x$ represents the input data, $W_{1}$ represents the weight matrix connecting the input layer to the hidden layer, $b$ denotes the bias vectors, and $W_{2}$ represents the weight matrix that links the hidden to output layer. $z_{1}$ and $z_{2}$ are binary vectors in the sampling process. MC-dropout approach is driven by the Bernoulli distributed random variables in the multiplying hidden activations. The random variables take the value of $1$ with probability parameter $p$ and 00 for other cases Gal2016 . This means the neuron is invalid if the corresponding binary variable takes 00 in a given input.
The $L_{2}$ regularization weighted by some weight decay gives a minimization objective shown as follows Huo2023 ; Wen2020 ,
$L_{\mathrm{dropout}}=\frac{1}{N}\sum_{n=1}^{N}||y_{n}-\hat{y}_{n}||_{2}^{2}+\lambda_{\mathrm{decay}}\sum_{i=1}^{L}(||W_{i}||_{2}^{2}+||b_{i}||_{2}^{2}).$ (2)
The first term denotes the Euclidean loss function used in the training and optimization process of neural network. Where $\hat{y}_{n}$ represents the output of a neural network with $L$ layers and $y_{n}$ denotes the target value. $N$ is the number of training data points. A regularization term is often added into the optimization process as shown in the second term. Here, the weight matrixes of dimensions $K_{i}\times{K_{i-1}}$ are represented by $W_{i}$ and the bias vectors of dimensions $K_{i}$ for each layer $i=1,2,\cdots,L$ are denoted by $b_{i}$. The quantity $\lambda_{\mathrm{decay}}$ represents the regularization parameter Huo2023 .
Figure 1: (Color online)Structure of the Monte Carlo dropout Bayesian neural network used in this work. The number of neurons in the input layer is $10$. The $3$ hidden layers are employed, and the number of neurons in the corresponding hidden layer $1$ and $3$ are $28$ and $71$, respectively. The hidden layer $2$, namely the dropout layer, has $300$ neurons. This diagrammatic sketch just shows the dropout hidden layer, in which the different markers denote the random variables take value of $1$ with probability parameter $p$ and 00 for other cases. The number of neurons in the output layer is $1$.
Shown in Refs. Zhang10662945 ; Hüllermeier2021 , the predicted errors of the observables are mainly originated from epistemic uncertainty and aleatoric uncertainty in using machine learning approaches. This means the chosen input structure parameter sets should capture the underlying physical mechanisms as much as possible. In this work, the experimental charge radii data of nuclei with $Z\geq 20$ and $A\geq 40$ are used to train the MC-dropout BNN PhysRevLett.126.032502 ; ANGELI201369 ; LI2021101440 . As mentioned in $Introduction$, the input structures for the MC-dropout BNN are consisted of the proton number ($Z$), mass number ($A$), isospin asymmetry degree ($I^{2}$), pairing effect ($\delta$), Casten factor ($P$), valence neutrons ($\nu_{n}$) and protons ($\nu_{p}$), quadrupole deformation ($\beta_{20}$), the hexadecapole deformation ($\beta_{40}$), and the profound shape-phase transition in 181,183,185Hg ($LI$). The ground state properties about quadrupole and hexadecapole deformation parameters are taken from Ref. MOLLER20161 .
The final input structures for the MC-dropout BNN approach are $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$, in which
$I^{2}$ $=$ $(\frac{N-Z}{A})^{2},$ (3)
$\delta$ $=$ $\frac{(-1)^{Z}+(-1)^{N}}{2},$ (4)
$P$ $=$ $\frac{\nu_{p}\nu_{n}}{\nu_{p}+\nu_{n}},$ (5)
$LI$ $=$ $\left\{\begin{array}[]{c c}1,&(Z,A)=(80,181),(80,183),(80,185)\\ 0,&\mathrm{else}\end{array}.\right.$ (8)
As mentioned before, the traditional Casten factor $P$ is invalid along the proton magic numbers $Z=28$, $50$, $82$ isotopic chains. To avoid this puzzle, the modified Casten factor $P^{*}$ is proposed as follows,
$P^{*}$ $=$ $\ln\left(\frac{e^{\nu_{p}}e^{\nu_{n}}}{{\nu_{p}}+{\nu_{n}}}\right).$ (9)
The input parameter sets which includes the traditional Casten factor $P$ cannot describe the shell quenching phenomena around the neutron magic numbers. In contrast to the traditional Casten factor $P$, the modified Casten factor $P^{*}$ can give more high accuracy in the training and validation sets, and the shell closure effect in charge radii can be reproduced well. More details can be found in the following $Section$. For the modified Casten factor $P^{*}$, the input structures can be redefined as $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P^{*}$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$. Meanwhile, if the proton numbers and neutron numbers simultaneously occupy the fully filled shells, such as $(Z=28,N=28)$, $(Z=28,N=50)$, $(Z=50,N=50)$, $(Z=50,N=82)$, and $(Z=82,N=126)$ cases, two kinds of Casten factors are vanished totally. In addition, the np formula Nerlo-Pomorska:1994dhg can also be chosen as our theoretical model to be refined by MC-dropout BNN:
$R_{\mathrm{np}}(Z,A)=r_{0}A^{1/3}\left(1-b\frac{N-Z}{A}+\frac{c}{A}\right),$ (10)
where $r_{0}=0.966$ fm, $b=0.182$, and $c=1.652$ Bayram:2013jua .
III Results and Discussion
For the convenience of discussion, we use “$P^{*}$” to denote the model combining NP formula and the MC-dropout BNN with input neurons $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P^{*}$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$. We also give the results obtained by the MC-dropout BNN with input neurons $z=\{Z,$ $A$, $I^{2}$, $\delta$, $P$, $\nu_{n}$, $\nu_{p}$, $\beta_{20}$, $\beta_{40}$, $LI\}$, denoted by “$P$” for comparison. Here, we should be mentioned that the difference between $P^{*}$ and $P$ just distinguishes the local variations in charge radii along $Z=28$, $50$, $82$ isotopic families. In the training set, we use the $815$ experimental data given in Ref. ANGELI201369 . The more recent experimental data PhysRevLett.126.032502 ; LI2021101440 , containing 118 data for nuclei with $Z\geq 20$ and $A\geq 40$, are used as the validation set to test the predictive power of our model. In contrast to Refs. PhysRevC.105.014308 ; DONG2023137726 , the training data sets are reduced owing to the whole data of charge radii along nickel isotopic chain are transformed into the validation data set.
To further quantify the predictive power, the rms deviation between the results obtained by the MC-dropout BNN model and the corresponding experimental data are calculated as follows,
$\sigma^{(\mathrm{T},\mathrm{V},\mathrm{TV})}=\sqrt{\frac{1}{N_{(\mathrm{T},\mathrm{V},\mathrm{TV})}}\sum_{i=1}^{N_{(\mathrm{T},\mathrm{V},\mathrm{TV})}}\left(R_{\mathrm{ch},i}^{\mathrm{pred}.}-R_{\mathrm{ch},i}^{\mathrm{exp}.}\right)^{2}}.$ (11)
where $N_{\mathrm{T}}$, $N_{\mathrm{V}}$, and $N_{\mathrm{TV}}$ are the numbers of charge radii contained in the training (T), validation (V), and entire (TV) data sets, and the subscript $i$ represents the $i$th nucleus in the given data sets.
As shown in Ref. PhysRevC.105.014308 , the predictive accuracy in describing the rms charge radii has been achieved to be $\sigma^{\mathrm{T}}=0.0143$ fm, and $\sigma^{\mathrm{TV}}=0.0149$ fm. Then the revised version has been proposed to further reduce the quantitative uncertainty DONG2023137726 , which leads to the $\sigma^{\mathrm{T}}=0.0140$ fm and $\sigma^{\mathrm{TV}}=0.0139$ fm. For these studies, the Casten factor associated to the valence neutrons and protons is regarded as input parameter. However, the quadrupole deformation has not been taken into account in the validated protocol. As demonstrated in Ref. PhysRevC.88.011301 , the rms deviation in charge radii can be reduced if the shape deformation is incorporated into the calibrated protocol. In recent study, the input sets $\{Z,A,\delta,\nu_{p},\nu_{n},\beta_{20}\}$ have been used in the training and validation data set, the rms deviation has been fallen into $\sigma^{\mathrm{T}}=0.0106$ fm and $\sigma^{\mathrm{TV}}=0.0130$ fm Zhang10662945 .
Table 1: Root-mean-square (rms) deviations of charge radii predicted by the $P$ and $P^{*}$ models.
Model $\sigma^{(\mathrm{T})}$ $\sigma^{(\mathrm{V})}$ $\sigma^{(\mathrm{TV})}$
P 0.0124 fm 0.0147 fm 0.0128 fm
P 0.0084 fm 0.0124 fm 0.0092 fm
In Table 1, the $\sigma^{\mathrm{T}}$, $\sigma^{\mathrm{V}}$, and $\sigma^{\mathrm{TV}}$ values derived from the Eq. (11) are shown, respectively. In our calculations, the valence neutrons, valence protons, Casten factor, and the shape-phase deformation in the limited region are simultaneously taken into account during the simulated process. Meanwhile, the higher deformation parameter $\beta_{40}$ is also included. The deduced value of $\sigma^{\mathrm{TV}}$ deriving from the $P$ model is equivalent to Ref. Zhang10662945 , but less than the results shown in Refs. PhysRevC.105.014308 ; DONG2023137726 . As mentioned above, the over-fitting puzzle can be encountered in the traditional Casten factor $P$ when meets the isotopic chains with proton numbers $Z=28$, $50$, $82$. The $\sigma^{\mathrm{TV}}$ value obtained by $P^{*}$ model can be further reduced to be $0.0094$ fm. This is accord with Ref. DONG2023137726 where the modified rms deviation has been used to measure the quantified uncertainty. In a word, the rms deviations in the training and validation data sets are also further reduced with respect to those investigations in the existing literature. It is therefore reasonable to demonstrate the predictive power of the $P^{*}$ model by modifying the traditional Casten factor $P$.
Significant improvement has been achieved by the $P^{*}$ model in both the interpolation and the extrapolation predictions of nuclear charge radii in comparison with the $P$ model. To visualize the details about the training and validation of the MC-dropout BNN approach, as shown in Fig. 2, the difference between the calibrated rms charge radii with $P^{*}$ model and the corresponding experimental data are depicted in nuclear chart. From this figure, it can be found that most of the experimental data can be reproduced well. In some specific regions, the predicted values are heavily deviated from the experimental data. As is well known, the inverse OES or the weakened OES can be evidently observed around Eu ($Z=63$Alkhazov1990 and At($Z=85$PhysRevC.99.054317 isotopes. In our calibrated protocol, the inverse odd-even staggering behavior has not been considered properly. This seems to lead the larger deviations between the predicted rms charge radii and the corresponding experimental data. Meanwhile, the larger deviations can also be encountered along manganese isotopes. This means that more underlying mechanisms should be covered through the input structures.
Figure 2: (Color online)Nuclear chart displaying the absolute differences (in units of fm) between the MC-Dropout BNN predicted rms charge radii and the experimental data in the entire set. The results are obtained from the training and validation procedures, respectively.
Figure 3: (Color online)Root-mean-square charge radii of calcium (a), nickel (b), tin (c), and lead (d) isotopes simulated by the $P^{*}$ and $P$ models, in comparison with the experimental data ANGELI201369 ; LI2021101440 . The results obtained by the np formula ($R_{\mathrm{np}}$) are also drawn for making a guide eye. The data covered by gray area are not contained in the training set, namely the pure predictions. The neutron magic numbers $N=20$, 28, 50, 82, and 126 are emphasized by the vertical lines.
Shell closure effects are naturally observed in charge radii throughout the whole nuclear chart ANGELI201369 ; LI2021101440 ; PhysRevLett.128.022502 ; PhysRevLett.129.132501 ; PhysRevLett.131.102501 . This attributes to the rather small isospin dependence of the spin-orbit interactions PhysRevLett.74.3744 . The relativistic mean field theory Geng:2003pk ; PhysRevC.102.024307 ; PhysRevC.109.064302 and the Fayans energy density functional model PhysRevC.95.064328 ; PhysRevLett.122.192502 can reproduce the shell quenching phenomena in nuclear charge radii well. However, as demonstrated in Ref. PhysRevLett.128.022502 , the Skyrme density functional theory cannot reproduce the shell closure effects. This provides an potential signature to identify the predictive power of nuclear models in describing the local variations of charge radii. In Fig. 3, the rms charge radii of calcium (a), nickel (b), tin (c), and lead (d) isotopes are depicted with both of the $P^{*}$ and $P$ models, respectively. As mentioned above, MC-dropout BNN model can provide the quantified uncertainty in the validated process. Therefore the corresponding error bars are also shown in Fig. 3. For calcium isotopes, $P^{*}$ and $P$ models give the similar trend of changes of charge radii. The slight deviations can be found at the neutron-deficient 34Ca and neutron-rich 54,55Ca isotopes. Meanwhile, charge radii of 39,43Ca isotopes can be reproduced well by the $P^{*}$ model in comparing with the $P$ model. Likewise, the systematical evolution of charge radii along tin isotopes can also be reproduced well by both of the $P^{*}$ and $P$ models as shown in Fig. 3 (c).
In Fig. 3 (b), it is obviously found that the $N=28$ shell closure effect cannot be reproduced well by the $P$ model. Besides, the charge radii of 54-56,58Ni isotopes are heavily overestimated. The same scenarios can also be found in the 65-68Ni isotopes, but with slight deviation. Meanwhile, charge radius of 70Ni obtained by the $P$ model is slightly underestimated with respect to $P^{*}$ model. Toward proton-rich regions, the pronounced deviations can be obviously exhibited between the predictive results obtained by the $P^{*}$ and $P$ models. These larger deviations can also be encountered in the neutron-rich lead isotopes as shown in Fig. 3 (d). Here, it should be mentioned that the rapid increase of charge radii can be reproduced well by the $P^{*}$ model, but the $P$ model fails to follow this trend beyond the neutron number $N=126$. All of these results suggest that $P^{*}$ model can provide a powerful prediction of nuclear charge radii. Therefore, the extrapolated simulation with high accuracy is necessary. In Table. 2, charge radii of calcium, nickel, tin, and lead isotopes are shown explicitly. These results are encouraged to provide an theoretical guidance for the experimental researches.
Table 2: The root-mean-square charge radii predicted by the $P^{*}$ and $P$ models, respectively.
       Nucleus        $P^{*}$ model        $P$ model
       34Ca        3.4059(56)        3.3920(44)
       35Ca        3.4156(53)        3.4193(37)
       53Ca        3.5721(40)        3.5800(62)
       54Ca        3.5956(48)        3.6089(65)
       55Ca        3.6046(47)        3.6267(72)
       51Ni        3.7470(40)        3.7687(33)
       52Ni        3.7517(40)        3.7682(35)
       53Ni        3.7448(35)        3.7581(35)
       57Ni        3.7412(56)        3.7686(38)
       69Ni        3.8976(32)        3.8942(36)
       71Ni        3.9066(28)        3.9019(32)
       72Ni        3.9128(27)        3.9105(33)
       105Sn        4.5303(40)        4.5348(51)
       106Sn        4.5414(36)        4.5446(47)
       107Sn        4.5517(33)        4.5519(46)
       133Sn        4.7183(38)        4.7162(52)
       135Sn        4.7375(36)        4.7355(45)
       136Sn        4.7471(39)        4.7450(42)
       137Sn        4.7545(40)        4.7527(42)
       179Pb        5.3638(35)        5.3688(41)
       180Pb        5.3718(34)        5.3740(42)
       181Pb        5.3787(33)        5.3785(42)
       213Pb        5.5441(62)        5.5326(44)
       215Pb        5.5642(85)        5.5426(44)
       216Pb        5.5744(99)        5.5479(44)
       217Pb        5.5808(111)        5.5525(44)
The input parameters which include the modified Casten factor $P^{*}$ can describe the systematical evolution of nuclear charge radii well. This can be easily found from the rms deviation of the training and validation data sets. Furthermore, as shown in Fig. 3, the shrunk phenomena in charge radii can be reproduced well with the $P^{*}$ model. Actually, Casten factor represents the neutron-proton correlation in describing the trend of changes of nuclear charge radii. As demonstrated in Refs. PhysRevC.53.1599 ; MILLER2019360 ; COSYN2021136526 , the neutron-proton correlation has an influence on determining the nuclear charge radii. This can also be found from Refs. PhysRevC.102.024307 ; PhysRevC.109.064302 where the neutron-proton correlation derived from the difference of the fractions of Cooper pairs between neutrons and protons can reproduce the shell quenching in charge radii as well. All of these suggest that more appropriate isospin interactions should be covered in describing the nuclear charge radii. Especially, the difference of charge radii can be used to ascertain the nuclear symmetry energy through the mirror-pair nuclei. Thus more available charge radii data are required in theoretical study.
IV Summary and conclusions
More input parameter sets physically motivated, such as proton number, mass number, isospin-asymmetry degree, pairing effect, shell closure effect, valence neutrons and protons, Casten factor $P$ or $P^{*}$, local shape-staggering phenomena, quadrupole deformation $\beta_{20}$, and higher order $\beta_{40}$ deformation, have been taken into account in the evaluated process. Significant improvements can be achieved in training and validating data sets. Especially, based on the $P^{*}$ model, the rms deviations for the training and validation stage fall into $\sigma^{\mathrm{T}}=0.0084$ fm and $\sigma^{\mathrm{V}}=0.0124$ fm, and $\sigma^{\mathrm{TV}}=0.0092$ fm for entire set. $P^{*}$ model can reproduce the shell quenching phenomena and the rapid increase of charge radii across neutron magic numbers. This suggests that $P^{*}$ model persists the predictive power when the theoretical uncertainties are taken into account. Even for nuclei located far from those contained in the training set, $P^{*}$ model can still make available predictions.
This work presents the potential approach of Bayesian neural network in exploring nuclear charge radii. We should mention that physically motivated features play an indispensable role in extrapolating cases far away from the $\beta$-stability line. Although the Monte Carlo dropout Bayesian neural network provides rather small root-mean-square deviation in simulating the theoretical results, the larger deviation can also be encountered in some local regions. This means the underlying physical mechanism cannot be captured adequately in the training and validation data sets. In addition, the number of training data sets has an influence on describing and predicting the nuclear charge radii. Therefore, more complex input structures should be taken into account properly in the proceeding research.
Acknowledgements
This work was supported by the Open Project of Guangxi Key Laboratory of Nuclear Physics and Nuclear Technology, No. NLK2023-05 and the Central Government Guidance Funds for Local Scientific and Technological Development, China (No. Guike ZY22096024).